### v1

# Privacy-Preserving Resilient Vector Consensus

Abstract
This paper studies privacy-preserving resilient vector consensus in multi-agent systems against faulty agents, where normal agents can achieve consensus within the convex hull of their initial states while protecting state vectors from being disclosed. Specifically, we consider a modification of an existing algorithm known as Approximate Distributed Robust Convergence Using Centerpoints (ADRC), i.e., Privacy-Preserving ADRC (PP-ADRC). Under PP-ADRC, each normal agent introduces multivariate Gaussian noise to its state during each iteration. We first provide sufficient conditions to ensure that all normal agents’ states can achieve mean square convergence under PP-ADRC. Then, we analyze convergence accuracy from two perspectives, i.e., the Mahalanobis distance of the final value from its expectation and the Hausdorff distance based alteration of the convex hull caused by noise when only partial dimensions are added with noise. Then, we employ concentrated geo-privacy to characterize privacy preservation and conduct a thorough comparison with differential privacy. Finally, numerical simulations demonstrate the theoretical results.
{IEEEkeywords}
Multi-agent systems, resilient vector consensus, geo-privacy
1 Introduction
[…]
Enabling multi-agent systems to collaborate effectively in solving complex problems or accomplishing tasks has garnered significant interest in fields such as aviation, robotics, and others [1, 2, 3]. In many application scenarios, such as distributed machine learning, multi-robot systems, and platoon, vector consensus is frequently used, aiming to make all agents’ state vectors reach agreement under a predefined rule. However, in multi-agent systems, there may be faulty or adversarial agents [^1] In the subsequent text, both are collectively referred to as faulty agents. against the rule to break the consensus of the system or drive the consensus to an unsafe value. To maintain the safe consensus of normal agents, the concept of resilient vector consensus has been developed. Resilient vector consensus means that all normal agents can achieve vector consensus where the final value lies within the convex hull of the initial states of normal agents despite the presence of faulty agents. Under resilient vector consensus, many distributed systems can work correctly under complex and changing circumstances. Therefore, resilient vector consensus is a very important concept in distributed fields.
In most consensus/resilient algorithms, agents directly send their states to their neighbors. However, if the transmitted states are intercepted by a malicious entity, this could lead to the disclosure of agents’ valuable information, resulting in a privacy leakage. For example, in distributed machine learning applications like the financial and medical sectors, the exchange of grid information during the training process may lead to the leakage of sensitive data from training sets. The sensitive data may include bank customers’ account balances, transaction records, as well as patients’ diagnosis results and medication histories. Such disclosure can potentially result in financial fraud and serious breaches of patient privacy, indicating that privacy preservation is a crucial task in the resilient vector consensus.
Due to the scarcity of communication and computation resources in distributed systems and the high demand of encryption algorithms on these resources, differential privacy has become the primary choice for privacy protection in consensus algorithms. Typically, in the resilient vector consensus, we aim to protect the initial state vectors of the normal agents without the existence of a trusted control center, where local differential privacy (local-DP) [4] is suitable to be utilized. Different from centralized differential privacy (central-DP), local-DP protects information during transmission from being intercepted, unrelated to differential attacks, and thus does not involve the concept of neighboring inputs. Local-DP also ensures that any pair of inputs generate similar outputs. However, the initial state range for each agent can be very large, and requiring pairs of distant initial states to produce similar outputs would reduce the system’s utility [^2] Utility refers to the extent to which a system can still perform its original functions after the application of privacy-preserving mechanisms.. This motivates us to use the concept of geo-privacy, characterizing the Euclidean distance between agents as one of the metrics for privacy protection. Specifically, we propose to use the concept of concentrated geo-privacy [5], which is a generalization of differential privacy, to describe the privacy preservation of resilient vector consensus. Apart from a more precise description of privacy preservation, it also offers advantages such as Gaussian mechanism and advanced composition.
Much attention has been paid to resilient consensus algorithms, which can be roughly classified into two categories. One is to detect and preclude faulty agents[6, 7, 8], while the other one is to try to find a safe value despite the presence of faulty ones[9, 10, 11, 12]. In this paper, we focus on the second one as it has more analytical results and related works regarding resilient scalar/vector consensus and privacy-preserving resilient scalar consensus are provided below.
Resilient scalar consensus is designed for the normal agents to converge to a common state that lies between the maximum and minimum values of normal agents’ initial states. There are two frequently used ways: One way is to use “median”. In [13], Zhang et al. propose a median consensus algorithm, where each normal agent updates by using its own value and the median of the states from its neighborhood. As for the other way, most of the algorithms rely on the strategy of simply disregarding suspicious values. For example, in a series of algorithms named Mean-Subsequence-Reduced (MSR) algorithm [9, 10], the main idea is to discard a fixed number of largest and smallest values in each normal agent’s neighborhood. Conversely, in the modification of Weighted MSR (W-MSR) algorithm [11, 12], each normal agent only discards values that are larger or smaller than its own to maintain its state.
Resilient vector consensus is ensured by that each normal agent must seek a point which is always located in the convex hull of its normal neighboring state vectors at each iteration. The existing work can be classified into three categories. The first category utilizes Tverberg partitions to compute Tverberg points. Specifically, it includes Byzantine vector consensus algorithm [14] and Approximate Distributed Robust Convergence algorithm [2]. The second category calculates the intersection of multiple convex hulls. This involves Algorithm $1$ proposed in [15] and the multidimensional approximate agreement algorithm described in [16]. The algorithm of the last category utilizes the concept of “centerpoint”, exemplified by “Approximate Distributed Robust Convergence Using Centerpoint” (ADRC) algorithm [17]. However, these algorithms are rather time-consuming, and most of them require approximation algorithms, which consequently reduce fault tolerance. Notably, among these algorithms, the ADRC [17] stands out with its better tolerance for the mature approximation algorithm. In each iteration, every normal agent computes the centerpoint (an extension of the median in higher dimensions) of its neighborhood and adjusts its position accordingly. Resilient vector consensus can be achieved as long as the communication topology and the number of faulty agents within the neighborhood meet certain requirements.
Privacy-preserving resilient scalar consensus can be achieved by two general approaches, i.e., cryptography-based methods [18, 19] and noise-adding-based methods [18, 20]. Although cryptography-based algorithms have higher convergence accuracy compared to noise-adding-based ones, encryption, information exchange, and decryption are time-consuming processes and thus significantly decrease efficiency. Besides, the noise-adding-based ones are typically more flexible, as it allows for the adjustment of noise intensity and distribution based on specific application scenarios and privacy requirements. In [20], Fiore et al. proposed a Differentially Private MSR (DP-MSR) algorithm, adding decaying, zero-mean Laplace noise to the scalar states of normal agents. They also analyzed resilient scalar consensus in the sense of probability and differential privacy.
Note that privacy-preserving resilient vector consensus is still an open issue. One simple way is to conduct a privacy-preserving resilient scalar consensus algorithm $d$ times, where $d$ is the dimension of states. However, this may cause the final value to fall outside the convex hull of the initial states of the normal agents [17]. Additionally, analyzing the performance of privacy-preserving resilient vector consensus by noise adding involves several key challenges: (i) The noise added makes the convergence non-deterministic and random, and the existing definition of resilient vector consensus is inapplicable. (ii) Due to the noise and resilience, we cannot derive the analytical solution of the final value, making it difficult to describe the accuracy of convergence. Additionally, high-dimensional convex hulls are inherently difficult to characterize and require specific values to obtain. After adding noise, the vertices and boundaries of the convex hull change randomly, making it challenging to characterize the convex hull after multiple iterations. (iii) At each iteration, each agent transmits and updates state information and the specific distribution of local state at each iteration is non-deterministic. We need to analyze the extent to which this information leaks the initial state, which is also a very difficult task.
To solve the above challenges, we consider a modification of ADRC, named Privacy-Preserving ADRC (PP-ADRC), and provide rigorous theoretical analysis for convergence and privacy performance. Compared to our conference version [21], we use Gaussian noise instead of Laplace noise and conduct a more rigorous performance analysis, Meanwhile, we add the final value distribution analysis by characterizing the deviation of the final value from the expectation. Moreover, we use CGP instead of DP to characterize the privacy-preservation and compare it with DP comprehensively. We also add more simulations under a $3-$dimensional case to illustrate our theoretical results. The contributions of this paper are given as follows.
  • We consider a modification of ADRC, named Privacy-Preserving ADRC (PP-ADRC), where each agent adds Gaussian noise to the local state vector for local interaction. We show sufficient conditions to ensure a resilient consensus of expectation.
  • We analyze the final value from two perspectives. First, we utilize the Mahalanobis distance to analyze the residuals of the final value from the expected one given a specified probability upper bound. Secondly, we employ the Hausdorff distance to assess the change between the convex hulls with and without noise with a specified probability upper bound, where only partial fixed dimensions are added with noise for each iteration.
  • We employ $\rho$-concentrated geo-privacy to characterize privacy preservation without detailed distributions of the outputs. Through a detailed comparison with differential privacy, we demonstrate that $\rho-$concentrated geo-privacy can provide advantages for resilient vector consensus.
The organization of this work is as follows. We introduce preliminaries and problem formulation in Sections 2 and 3, respectively. Then, we analyze the convergence condition and final value in Section 4. The privacy analysis is presented in Section 5. Simulation results are presented in Section 6. Finally, conclusions and avenues for future research are outlined in Section 7.
2 PRELIMINARIES
In this section, we introduce some preliminaries on the basic notations, reachable graph sequence, and privacy notions.
Basic Notations: Throughout this paper, we denote by $\mathbb{R}^{d}$ the $d$-dimensional Euclidean space, by $\mathbb{R}^{m\times d}$ the space of all $m\times d$-dimensional matrices, and by $(\mathbb{R}^{m\times d})^{\mathbb{N}}$ the space of matrix valued sequences in $\mathbb{R}^{m\times d}$, where $\mathbb{N}$ denotes the set of natural numbers. The natural logarithm is denoted by $\log$. Let $\mathbf{A}$ be a matrix. Then, we denote by $[\mathbf{A}]_{i},[\mathbf{A}]^{j},[\mathbf{A}]_{ij},\mathbf{A}^{\top}$ the $i$th row, the $j$th column, the $(i,j)$th element, and the transpose of matrix $\mathbf{A}$. We then consider two matrices $\mathbf{A}$ and $\mathbf{B}$, and if $[\mathbf{A}]_{ij}\leq[\mathbf{B}]_{ij}$ holds for any pair of $i,j$, we say $\mathbf{A}\leq\mathbf{B}$. For a vector $x\in\mathbb{R}^{d}$, we denote by $x_{\operatorname{max}}$ and $x_{\operatorname{min}}$ maximum element and minimum element of it, respectively. Then, we consider a matrix consisting of $n$ row vectors $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]^{\top}$, where $x_{i}$ denotes the $i$th vector and $x_{i,k}$ stands for the $k$th element of the vector $x_{i}$. We denote by $\mathbf{1}$ a column vector of ones, and by $I_{d}$ the $d-$dimensional identity matrix. A matrix is row-stochastic if all its elements are non-negative and each row sums to $1$. For a given point set $C\subseteq\mathbb{R}^{d}$, we denote by $\text{conv}(C)$ the convex hull of the point set $C$, by $|C|$ the cardinality of $C$. We consider two functions $f(x)$ and $g(x)$, and $f(x)=\Omega(g(x))$ means that $f(x)\geq g(x)$. For two different vectors $x,x^{\prime}\in\mathbb{R}^{d}$, we represent by $\mathrm{dist}(x,x^{\prime})=||x-x^{\prime}||_{2}$ the Euclidean distance between $x$ and $x^{\prime}$. For two $n$-tuples of vectors $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]^{\top}$ and $\mathbf{x}^{\prime}=[x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{n}^{\prime}]^{\top}$, $\mathrm{dist}(\mathbf{x},\mathbf{x}^{\prime})$ stands for the maximum distance among all the vectors, which is $\max\limits_{i}\operatorname{dist}(x_{i},x_{i}^{\prime})$. In probability theory, given a random variable $z\in\mathbb{R}$, we denote by $\mathbb{E}(z)$ and $\operatorname{var}(z)$ the expectation and the variance of $z$, respectively. As for a random vector $z=[z_{1},\ldots,z_{d}]^{\top}\in\mathbb{R}^{d}$, the expectation of $z$ is $\mathbb{E}(z)=[\mathbb{E}(z_{1}),\ldots,\mathbb{E}(z_{d})]^{\top}$. For a topological space $\mathcal{T}$, we denote by $\mathcal{B}(\mathcal{T})$ the set of Borel subsets of $\mathcal{T}$, and $\mathrm{P}$ is the corresponding probability measure. For a zero mean $d-$dimensional Gaussian distribution, the probability density function is given by $G(x;\Sigma)=\frac{1}{\sqrt{(2\pi)^{d}|\Sigma|}}\mathrm{exp}\left(-\frac{1}{2}x^{\top}\Sigma^{-1}x\right)$.
Reachable Graph Sequence: We denote by $\mathcal{G}(t)=(\mathcal{V},\mathcal{E}(t)),\text{ }t=0,1,2,\ldots$ a time-varying directed graph, where $\mathcal{V}$ is the set of vertices, and $\mathcal{E}(t)\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. Then, we introduce two definitions of reachability with a relationship to a graph sequence below [2].
Definition 1
(Jointly Reachable Graph Sequence): Given $j\in\mathbb{N}$ and a finite time sequence $T_{j},T_{j}+1,\ldots,T_{j+1}-1$, the corresponding finite sequence of graphs $\mathcal{G}(T_{j}),\mathcal{G}(T_{j}+1),\ldots,\mathcal{G}(T_{j+1}-1)$ is said to be a jointly reachable graph sequence if in the union of graphs $\bigcup_{t=T_{j}}^{T_{j+1}-1}\mathcal{G}(t)=\left(\mathcal{V},\bigcup_{t=T_{j}}^{T_{j+1}-1}\mathcal{E}(t)\right)$, there is a vertex $v\in\mathcal{V}$ such that $\forall v^{\prime}\neq v$, we can find a path from $v^{\prime}$ to $v$ in the union of graphs.
Definition 2
(Repeatedly Reachable Graph Sequence): An infinite sequence of graphs $\mathcal{G}(0),\mathcal{G}(1),\ldots$ is said to be a repeatedly reachable graph sequence if there is an infinite time sequence, $T_{\mathrm{inf}}:0=T_{1}<T_{2}<\ldots$, such that for any $j\in\mathbb{N}$, the subsequence $\mathcal{G}(T_{j}),\mathcal{G}(T_{j}+1),\ldots,\mathcal{G}(T_{j+1}-1)$ is a jointly reachable graph sequence.
Privacy Notions: We first give two formal definitions of standard local differential privacy and then introduce CGP.
Definition 3
($\varepsilon-$Differential Privacy): Given spaces $U$, $V$, and $\varepsilon\in\mathbb{R}\geq 0$, a randomized function $M:U\rightarrow V$ is $\varepsilon-$differentially private, if for any pair of inputs $x$, $x^{\prime}\in U$ and any $S\subseteq V$, we have $\mathrm{P}\{M(x)\in S\}\leq e^{\varepsilon}\mathrm{P}\{M(x^{\prime})\in S\}.$
Definition 4
($(\varepsilon,\delta)-$Differential Privacy): Given spaces $U$, $V$, and $\varepsilon$, $\delta\in\mathbb{R}\geq 0$, a randomized function $M$ is $(\varepsilon,\delta)-$differentially private, if for any pair of inputs $x$, $x^{\prime}\in U$ and any $S\subseteq V$, $\mathrm{P}\{M(x)\in S\}\leq e^{\varepsilon}\mathrm{P}\{M(x^{\prime})\in S\}+\delta$ holds.
Note that in Definitions 3-4, the inputs $x$ and $x^{\prime}$ can be in the real domain that includes one or more scalars or vectors. The parameter $\varepsilon$ in both definitions characterizes the privacy protection strength, with smaller values of $\varepsilon$ providing stronger privacy protection. The unique difference between the two definitions is the relaxation term $\delta$, i.e., $(\varepsilon,\delta)$-DP means that $\varepsilon$-DP is achieved with a probability of at least $1-\delta$. It is noteworthy that the Gaussian mechanism can only achieve $(\varepsilon,\delta)$-DP.
Definitions 3-4 are quite strict for resilient vector consensus. As the range of possible initial states of an agent is typically large, ensuring two far apart initial states produce similar outputs would cause significant system utility reduction. To solve this issue, geo-privacy has been widely studied, which uses Euclidean distance as one metric for privacy protection. Supporting the Gaussian mechanism and better composition, CGP is selected here, which is established on Rényi divergence given below [22].
Definition 5
(Rényi Divergence): Given two distributions $F$ and $G$ on domain $\operatorname{dom}(x)$ with pdf $f(x)$ and $g(x)$, respectively, Rényi divergence of order $\alpha>1$ is defined as
$D_{\alpha}(F\|G)$ $=\frac{1}{\alpha-1}\log\left(\int_{\operatorname{dom}(x)}f(x)^{\alpha}g(x)^{1-\alpha}\mathrm{d}x\right)$ (1)
$=\frac{1}{\alpha-1}\log\left(\underset{x\thicksim f(x)}{\mathbb{E}}\bigg{[}\left(\frac{f(x)}{g(x)}\right)^{\alpha-1}\bigg{]}\right).$
Note that $D_{\alpha}(F\|G)$ is determined by the expectation of the $\alpha-1$ power of the ratio of two distributions at each point, quantifying the difference between two distributions. Meanwhile, larger $\alpha$ emphasize parts where the ratio is greater, while smaller values of $\alpha$ focus more on the overall average differences. The roles of logarithm and $\frac{1}{\alpha-1}$ are both intended to prevent numerical overflow in the results. Rényi divergence is monotonically non-decreasing with $\alpha$.
Definition 6
(Concentrated Geo-Privacy): Given spaces $U$, $V$, and $\rho\in\mathbb{R}\geq 0$, a randomized function $M:U\rightarrow V$ is said to satisfy $\rho-$concentrated geo-privacy, if for any inputs $x,x^{\prime}\in U$ and all $\alpha>1$, it holds that
$D_{\alpha}(M(x)\|M(x^{\prime}))\leq\alpha\rho\cdot\mathrm{dist}(x,x^{\prime})^{2}.$ (2)
In CGP, $\rho$ provides the upper bound on $D_{\alpha}(M(x)\|M(x^{\prime}))/\alpha/\mathrm{dist}(x,x^{\prime})^{2}$, characterizing the difference between the output distributions of $M$ for any pair of inputs, thereby quantifying the strength of privacy protection. A smaller $\rho$ indicates that $M$ is less sensitive to the inputs, thereby providing stronger privacy protection. Besides, the privacy should degrade gracefully as $\mathrm{dist}(x,x^{\prime})$ increases.
3 Problem Formulation
In this section, we first provide the network model of the multi-agent system with faulty agents and then propose the PP-ADRC algorithm. Finally, interesting problems are stated.
3.1 Network Model
We consider a multi-agent system with $n$ agents modeled by a time-varying directed graph $\mathcal{G}(t)=(\mathcal{V},\mathcal{E}(t)),\text{ }t=0,1,2,\ldots$, where $\mathcal{V}=\{1,2,\ldots,n\}$ is the set of agents, and $\mathcal{E}(t)\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. The set of in-neighbors of an agent $i$ is denoted by $\mathcal{N}_{i}^{\text{in}}(t)=\{j\in\mathcal{V}|(j,i)\in\mathcal{E}(t)\}$, while the set of out-neighbors is $\mathcal{N}_{i}^{\text{out}}(t)=\{j\in\mathcal{V}|(i,j)\in\mathcal{E}(t)\}$, and the corresponding out-degree is $d^{\text{out}}_{i}(t)=|\mathcal{N}_{i}^{\text{out}}(t)|$ with $d^{\text{out}}_{\text{max}}=\max\limits_{\{i,t\}}(d^{\text{out}}_{i}(t))$.
The system has two types of agents, i.e., normal agents and faulty agents. Each normal agent updates its state through local interaction based on predefined rules. Conversely, faulty agents can behave arbitrarily and unpredictably. They are classified into two categories: Byzantine agents and malicious agents. Byzantine agents send different arbitrary states to different neighbors, while malicious agents can only send the same state to all neighbors [11]. Malicious agents are actually a specific type of Byzantine agents. Therefore, for a more general case, we assume that the faulty agents in this paper are Byzantine agents. We denote by $\mathcal{F}\in\mathcal{V}$ the set of faulty agents, and by $f=|\mathcal{F}|$ the total number of faulty agents with $f\leq F$. For any $i\in\mathcal{F}$ and $j\in\mathcal{V}$, we denoted by $\hat{x}_{i}^{j}(t)$ the state $i$th faulty agent sent to $j$th agent at iteration $t$. Let $\overline{\mathcal{V}}=\mathcal{V}-\mathcal{F}$, and $\overline{n}=n-f$, where $\overline{\mathcal{V}}$ is the set and $\overline{n}$ is the number of all normal agents. Without loss of generality, we suppose $\overline{n}$ ones in front are the normal agents, meaning that $\overline{\mathcal{V}}=\{1,2,\ldots,\overline{n}\}$. Besides, we denote by a $\overline{n}\times d$ matrix $\overline{\mathbf{x}}(t)=[x_{1}(t),x_{2}(t),\ldots,x_{\overline{n}}(t)]^{\top}$ the states of all normal agents, where $x_{i}(t)$ denotes the state of a normal agent, and by $\overline{\mathbf{x}}_{0}$ the initial states of normal agents. We define the topology of normal agents by a time-varying directed graph $\overline{\mathcal{G}}(t)=(\overline{\mathcal{V}},\overline{\mathcal{E}}(t))$, where ¯$\overline{\mathcal{E}}(t)\subseteq{\overline{\mathcal{V}}}\times\overline{\mathcal{V}}$. For each normal agent $i\in\overline{\mathcal{V}}$, we denote by $\overline{\mathcal{N}}_{i}^{\text{in}}(t)$ and $n_{f_{i}}(t)$ its set of in-neighbors in $\overline{\mathcal{G}}(t)$ and the number of its faulty in-neighbors in $\mathcal{G}(t)$, respectively.
3.2 Algorithm Design
We consider a modification of the existing ADRC Using Centerpoint (ADRC) algorithm, i.e., Privacy-Preserving ADRC (PP-ADRC) algorithm. Its details are presented in the Algorithm 1. To protect privacy, zero-mean, decaying Gaussian noise is added to the state of each normal agent during the communication phase. In ADRC, the key to achieving resilience lies in calculating the centerpoint $s_{i}(t)$ during the calculation phase.
Algorithm 1 PP-ADRC
  Input: $\mathcal{G}(t)$, $\overline{\mathbf{x}}(t)$, $\Sigma(t)=\lambda^{2}\upsilon^{2t}I_{d}$, $\gamma_{i}(t)$, threshold $\mathrm{thre}$
  Output: final value $x(\infty)$
  Initialization: initialize states of normal agents $\overline{\mathbf{x}}_{0}$, $t=0$
  Iteration:
  for each normal agent $i\in\overline{\mathcal{V}}$ do
     Transmission phase:
     Add noise to its state, i.e.,
$y_{i}(t)=x_{i}(t)+\eta_{i}(t),$ (3)
where $\eta_{i}(t)\in\mathbb{R}^{d}$ is a zero-mean decaying $d$-dimensional Gaussian noise with covariance matrix $\Sigma(t)$.
     Transimit $y_{i}(t)$ to its out-neighbors.
     Calculation phase:
     Calculate centerpoint $s_{i}(t)$.
     Update phase:
     Update its state following:
$x_{i}(t+1)=\gamma_{i}(t)s_{i}(t)+(1-\gamma_{i}(t))x_{i}(t),$ (4)
where $10<\gamma_{l}\leq\gamma_{i}(t)\leq\gamma_{m}<$ 1 and $\gamma_{l}>1-\upsilon$.
  end for
  If $||x_{i}(t+1)-x_{i}(t)||_{2}<\mathrm{thre},\forall i\in\overline{\mathcal{V}}$, the iteration terminates.
For a set of $n$ points in general position in $\mathbb{R}^{d}$, there always exists a centerpoint $p$, which is guaranteed to lie within the convex hull of any subset containing more than $\frac{nd}{d+1}$ points from the given point set [23, 24]. Therefore, if the fraction of faulty agents in a normal agent’s neighborhood is lower than $\frac{1}{d+1}$, the centerpoint always lies in the convex hull formed by the normal agents’ state vectors. However, the fraction $\frac{1}{d+1}$ is just a theoretical value. When the dimensionality exceeds three, calculating a centerpoint becomes a coNP-Complete problem, and we can only use approximation algorithms. Therefore, in practice, if $d>3$, we can only achieve a fraction of $\Omega\left(\frac{1}{d^{2}}\right)$ [17].
It should be pointed out that we choose the noise with zero mean Gaussian distribution with covariance matrix $\Sigma(t)$ for two reasons: 1) The properties of zero-mean and decaying variance are the necessary conditions to reach resilient vector consensus; 2) Gaussian distribution, whose pdf is also proportional to the $2$-norm, facilitates our subsequent privacy analysis when we use the distance between agents (measured by the $2$-norm) as a parameter for evaluating privacy preservation. Additionally, Gaussian noise has a smaller variation in magnituede with dimension changes compared to Laplace noise.
3.3 Problem of Interests
Different from the resilient vector consensus without adding noise, the randomness of noise makes the result under PP-ADRC become non-deterministic. Thus, we need to characterize metrics for resilient vector consensus and privacy preservation after noise addition, which are provided below.
Definition 7
(Resilient Vector Consensus): The resilient vector consensus is achieved if state vectors of all normal agents satisfy the following two conditions:
  • Safety: For any normal agent $i\in\overline{\mathcal{V}}$ and any $t>0$, $\mathbb{E}[x_{i}(t)]$ must be in the convex hull formed by the initial states of the normal agents, i.e. $\mathbb{E}[x_{i}(t)]\in\operatorname{conv}\left(x_{1}(0),x_{2}(0),\ldots,x_{\overline{n}}(0)\right)$.
  • Agreement: For any pair of normal agents $i,j\in\overline{\mathcal{V}}$, there always holds $\lim\limits_{t\rightarrow\infty}\mathbb{E}[||x_{i}(t)-x_{j}(t)||_{2}^{2}]=0$.
To characterize the privacy preservation of the state vector sequences, we first denote by three matrices $\overline{\mathbf{x}}(t),\ \overline{\boldsymbol{\eta}}(t),\ \text{and}\ \overline{\mathbf{y}}(t)\in\mathbb{R}^{\overline{n}\times d}$ the states, noises, and noised states of all the agents at time $t$, respectively. Then, we define the sequences of matrices $\mathbf{X}=\{\overline{\mathbf{x}}(t)\}_{t=0}^{\infty},\mathbf{N}=\{\overline{\boldsymbol{\eta}}(t)\}_{t=0}^{\infty},\ \text{and}\ \mathbf{Y}=\{\overline{\mathbf{y}}(t)\}_{t=0}^{\infty}$ from $t=0$ to $\infty$, and they are in the sample space $\Omega=(\mathbb{R}^{\overline{n}\times d})^{\mathbb{N}}$. Given an initial state $\overline{\mathbf{x}}_{0}$, the sequences $\mathbf{X}$ and $\mathbf{Y}$ are uniquely determined by the noise sequence $\mathbf{N}$ and we define the corresponding function as $\mathbf{Y}_{\overline{\mathbf{x}}_{0}}(\mathbf{\mathbf{N}})=\mathbf{\mathbf{X}}$.
Definition 8
(CGP in Resilient Vector Consensus): Given any pair of initial states $\overline{\mathbf{x}}_{0}$, $\overline{\mathbf{x}}_{0}^{\prime}$, and $\rho>0$, the system is said to satisfy $\rho-$CGP if and only if for all $\alpha>1$
$D_{\alpha}(\mathbf{y}_{\overline{\mathbf{x}}_{0}}(\mathbf{N})\|\mathbf{y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N}))\leq\alpha\rho\cdot\mathrm{dist}(\overline{\mathbf{x}}_{0},\overline{\mathbf{x}}_{0}^{\prime})^{2},$ (5)
where $\mathbf{y}_{\overline{\mathbf{x}}_{0}}(\mathbf{N})$ and $\mathbf{y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N})$ are the pdfs of $\mathbf{Y}_{\overline{\mathbf{x}}_{0}}(\mathbf{N})$ and $\mathbf{Y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N})$, respectively.
After noise adding, convergence analysis and privacy quantization become challenging. For convergence analysis, we can only describe the situation in terms of expectation. But for the final value, although the expectation is ensured to lie within the convex hull of normal agents’ initial states, we cannot obtain an analytical solution. Specifically, the exact result of each convergence and its deviation from the expectation are unknown. Additionally, characterizing the change of the convex hull after adding noise is very challenging because the vertices and boundaries are constantly changing. Without analytical expressions, it is hard to quantitatively describe the changes in the convex hull. For privacy preservation, we add an infinite amount of noise, and the output is an infinite sequence of state matrices without an explicit distribution expression. Analyzing the Rényi divergence of such infinite matrix sequences is inherently complex and necessitates a specialized approach. Consequently, we are interested in the following questions of the system performing PP-ADRC:
  1. 1.
    Under what conditions can we demonstrate that resilient vector consensus can be achieved?
  2. 2.
    How can we determine the change in the convex hull of the initial states of normal agents after adding noise? Furthermore, although the expectation of the final value remains within the original convex hull, how can we quantify the deviation from it for each single final value?
  3. 3.
    How can we derive the upper bound of $\rho$-CGP without knowing the specific distribution of two infinite matrix sequences? What are the advantages compared to differential privacy?
4 Convergence Analysis
In this section, we provide sufficient conditions to ensure resilient vector consensus. Then, we analyze the convergence accuracy from two perspectives, i.e., the deviation of the final value from the expectation and the change of the convex hull.
4.1 Resilient Vector Consensus
We first formulate the state evolution of all normal agents as a linear time-varying (LTV) system by leveraging the properties of the centerpoint. Then, by sequentially multiplying the iterative formula of the LTV system from the initial time step, we can leverage repeated reachability and the characteristics of stochastic matrices to establish the resilient vector consensus of PP-ADRC [2].
Lemma 1
Under PP-ADRC, if
$n_{f_{i}}(t)<N_{f_{i}}(t)=\begin{cases}\frac{|\mathcal{N}_{i}^{\text{in}}(t)|}{d+1}\ &\text{if }d=2,3\\ \Omega(\frac{|\mathcal{N}_{i}^{\text{in}}(t)|}{d^{2}})\ &\text{if }d>3\end{cases}$ (6)
holds for any $t\geq 0$ and any $i\in\overline{\mathcal{V}}$, then we have
$\overline{\mathbf{x}}(t+1)=\mathbf{M}(t)\overline{\mathbf{x}}(t)+\mathbf{H}(t)\mathbf{v}(t),\ t=0,1,2,\ldots$ (7)
where $\mathbf{M}(t)\in\mathbb{R}^{\overline{n}\times\overline{n}}$ is a row-stochastic matrix with $[\mathbf{M}(t)]_{ii}=1-\gamma_{i}(t)$, $\mathbf{H}(t)\in\mathbb{R}^{\overline{n}\times\overline{n}}$ has zero diagonal entries ans differs from $\mathbf{M}(t)$ only in the diagonal elements, and $\mathbf{v}(t)=\left[\eta_{1}(t),\eta_{2}(t),\ldots,\eta_{\overline{n}}(t)\right]^{\top}$.
We denote by $\Psi_{t,t_{s}}$ the backward product of the sequence $\{\mathbf{M}(t),t\geq 0\}$, i.e., $\Psi_{t,t_{s}}:=\mathbf{M}(t-1)\dots\mathbf{M}(t_{s}),\mathrm{for}\text{ }t>t_{s}\geq 0,\text{ }\Psi_{t,t}:=I$. Then, we obtain
$\overline{\mathbf{x}}(t+1)=\Psi_{t+1,0}\overline{\mathbf{x}}(0)+\sum_{q=0}^{t}\Psi_{t+1,q+1}\mathbf{H}(q)\mathbf{v}(q).$ (8)
Theorem 1
Under PP-ADRC, if equation (6) holds and $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, then for any $i\in\overline{\mathcal{V}}$, there exists a random vector $x(\infty)=[x_{i,1}(\infty),\ldots,x_{i,d}(\infty)]^{\top}$ such that $\lim\limits_{t\rightarrow\infty}\mathbb{E}\left[||x_{i}(t)-x(\infty)||_{2}^{2}\right]=0$, and $\mathbb{E}[x_{i}(t)]\in\operatorname{conv}\left(x_{1}(0),x_{2}(0),\ldots,x_{\overline{n}}(0)\right)$, where $t=0,1,2,\ldots$.
Proof 4.2.
Here. we use ergodicity333 Ergodicity implies that for the backward product of a row-stochastic matrix sequence, as time approaches infinity, each row becomes identical. to show the convergence. With equation (8), by using the independence and zero-mean of noise, for any $i\in\overline{\mathcal{V}}$, we obtain
$\lim\limits_{t\rightarrow\infty}\mathbb{E}\left[||x_{i,k}(t+1)-x_{j,k}(t+1)||_{2}^{2}\right]=0,\ \text{and}$ (9)
$\lim\limits_{t\rightarrow\infty}\mathbb{E}\left[||x_{i,k}(t+1)-x_{i,k}(t)||_{2}^{2}\right]=0.$
Hence, there must exist a random variable $x_{i,k}(\infty)$ satisfying $\lim\limits_{t\rightarrow\infty}\mathbb{E}\left[||x_{i,k}(t)-x_{i,k}(\infty)||_{2}^{2}\right]=0$, i.e.,
$\lim\limits_{t\rightarrow\infty}\mathbb{E}\left[||x_{i}(t)-x(\infty)||_{2}^{2}\right]=0,\ \forall i\in\overline{\mathcal{V}}.$ (10)
Additionally, by the zero-mean and independence of noise, we obtain $\mathbb{E}\left[\overline{\mathbf{x}}(t+1)\right]=\mathbb{E}\left[\Psi_{t+1,0}\overline{\mathbf{x}}(0)\right]$. As $\Psi_{t+1,0}$ is row-stochastic, one derives $\mathbb{E}\left[x_{i}(t+1)\right]\in\operatorname{conv}\left(x_{1}(0),x_{2}(0),\ldots,x_{\overline{n}}(0)\right)$.
Remark 4.3.
It should be pointed out that equation (6) ensures “safety”, while repeated reachability ensures the ergodicity of backward product $\Psi_{t,t_{s}}$, thereby guaranteeing “agreement”. After adding noise, we can only guarantee the “mean square convergence” under expectation. As the goal is to characterize the effect of noise term by expectation, independence, and zero-mean are the crucial factors, rather than the specific distribution of the noise. This means that resilient vector consensus can be achieved for other noise distributions with independence and zero-mean properties.
4.2 Distribution of Final Value
From Theorem 1, although the expectation of the final value is in the convex hull of the normal agents’ initial states, we cannot guarantee that each single convergence result always lies in the convex hull. This uncertainty arises because the specific distribution of the final value remains unknown. Thus, we investigate the accuracy of each final value in terms of its residuals, i.e., the deviation from the expected value. Specifically, we propose to use the Mahalanobis distance [25] and the multivariate Chebyshev’s inequality [26] to analyze the residuals. The details are provided as follows.
Definition 4.4.
Given a random vector $z$ and its non-singular covariance matrix $\Sigma$, the Mahalanobis distance $D_{M}(z)$ from $z$ to its expectation $\mathbb{E}(z)$ is defined as
$D_{M}(z)=\sqrt{(z-\mathbb{E}(z))^{T}\Sigma^{-1}(z-\mathbb{E}(z))}.$ (11)
In resilient vector consensus, our primary concern is the convergence accuracy of the random vector $x(\infty)$. The commonly used Euclidean distance is not suitable in this scenario due to varying variances of each component in $x(\infty)$ and existing correlations among them. These factors introduce different measurement scales across dimensions. Therefore, we opt for the Mahalanobis distance to quantify the deviation of the final value from its expected value. This metric offers a comprehensive approach to distance measurement by incorporating variances and correlations between components. By mitigating the influence of disparate scales on distance calculation, the Mahalanobis distance provides a more precise evaluation.
Lemma 4.5.
For a random vector $z\in\mathbb{R}^{d}$ with non-singular covariance matrix $\Sigma$, we have
$\mathrm{P}\left\{[z-\mathbb{E}(z)]^{\top}\Sigma^{-1}[z-\mathbb{E}(z)]\geq\chi\right\}\leq\frac{d}{\chi},\quad\forall\chi>0.$ (12)
We then show that covariance matrix $x(\infty)$ is non-singular. Combining Definition 4.4 and Lemma 4.5, we can derive the result below.
Theorem 4.6.
Under PP-ADRC, if equation (6) holds, $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, and the covariance matrix of $x(\infty)$ is non-singular, then
$\mathrm{P}\left\{[D_{M}(x(\infty))]^{2}\leq\chi\right\}\geq 1-\frac{d}{\chi},\quad\forall\chi>0.$
Remark 4.7.
The points sharing the same Mahalanobis distance from the expected value define a hyperspheroid centering at the expectation of the final value. Therefore, Theorem 4.6 encapsulates the final value within a hyperspheroid centered at a point (the expectation) inside the convex hull. In fact, if we know the position of the expectation, we can determine the probability that the final value lies within the convex hull more specifically. Additionally, note that the principal axis parameters of the hyperspheroid are determined by the eigenvalues of the covariance matrix and $\chi$ together. Thus, we can show that the stronger the noises are, the bigger the hyperspheroid is, implying a lower convergence accuracy. Particularly, the volume of the hyperspheroid is $V=V_{d}\left[\sqrt{|\Sigma|\chi}\right]^{d}$, where $V_{d}$ characterizes the volume of a d-dimensional unit hypersphere [27]. Meanwhile, we can write the covariance matrix $\Sigma$ as
$\left.[\Sigma]_{kk^{\prime}}=\left\{\begin{array}[]{ll}\operatorname{var}(x_{i,k}(\infty))&\text{if }k=k^{\prime}\\ \tau_{k,k^{\prime}}\sqrt{\operatorname{var}(x_{i,k}(\infty))\operatorname{var}(x_{i,k^{\prime}}(\infty))}&\text{if }k\neq k^{\prime},\end{array}\right.\right.$ (13)
where $\tau_{k,k^{\prime}}$ is the correlation coefficient. Consequently, we can derive the determinant of $\Sigma$, i.e., $|\Sigma|=h(\tau_{1,2},\ldots,\tau_{d-1,d})\operatorname{var}(x_{i,1}(\infty))\times\cdots\times\operatorname{var}(x_{i,d}(\infty))$, where $h(\cdot)$ represents a polynomial. It means that the volume is directly proportional to the variances.
For the case where the covariance matrix of $x(\infty)$ is singular, we can obtain the following results [28].
Lemma 4.8.
For a random vector $z\in\mathbb{R}^{d}$, if its covariance matrix $\Sigma$ is singular, then there exists at least one non-zero row vector $\zeta$ such that $\operatorname{var}(\zeta z)=0$, which means $\zeta z$ is a constant.
Remark 4.9.
Lemma 4.8 implies that certain components of the random variable can be expressed as linear polynomials of the remaining components. Identifying these extraneous components allows us to obtain a random vector $Z^{\prime}\in\mathbb{R}^{d^{\prime}}$ with a non-singular covariance matrix $\Sigma^{\prime}$. We consider the case where the covariance matrix of $x(\infty)$ is singular and without loss of generality, the latter $d-d^{\prime}$ components of $x(\infty)$ are supposed to be extraneous. It means that Theorem 5.20 still holds for the remaining $d^{\prime}$ components, where the result simply degenerates into a $d^{\prime}$-dimensional ellipsoid (or a line if $d^{\prime}=1$).
4.3 Convex Hull Change
In contrast to scalar consensus, the high-dimensional convex hull is defined by numerous vertices, dimensions, and coefficients. Moreover, the stochastic nature of noise can significantly alter the convex hull of typical agents, making it challenging to characterize. Furthermore, following a single iteration, the new convex hull may exhibit no overlap with the original configuration.
To solve this issue, we consider a special case that only partial dimensions need to be protected. Consequently, the convex hull of dimensions without adding noise retains its original form. Our focus shifts to analyzing the transformation of the convex hull in the remaining dimensions of the state. It is worth noting that analyzing high-dimensional convex hulls can be challenging, while the $1$-dimensional final value $x_{i,k}(\infty)$ can be assessed by examining its expectation and variance. Hence, we start convex hull change analysis from the perspective of one dimension.
We first evaluate the expectation and variance of $x_{i,k}(\infty)$. Then, we determine the one-dimensional convergence accuracy using Chebyshev’s inequality. Finally, we quantize the change of the convex hull caused by the Gaussian noise and its associated probability. The properties of the random variable $x_{i,k}(\infty)$ are given below.
Lemma 4.10.
Under PP-ADRC, if equation (6) holds, and $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, then
$\mathbb{E}\left[x_{i,k}(\infty)\right]=\psi[\overline{\mathbf{x}}_{0}]^{k},\operatorname{var}(x_{i,k}(\infty))\leq\frac{\lambda^{2}}{1-\upsilon^{2}},1\leq k\leq d,$
where $\psi$ is a $1\times\overline{n}$ non-negative row vector satisfying $\psi\mathbf{1}=1$.
Proof 4.11.
Similar to that in [20], using the independence and zero-mean of noise, we can obtain
$\mathbb{E}\left[x_{i,k}(\infty)\right]=\lim\limits_{t\rightarrow\infty}\mathbb{E}\left[\Psi_{t+1,0}[\overline{\mathbf{x}}_{0}]^{k}\right]=\psi[\overline{\mathbf{x}}_{0}]^{k}.$
Then for the variance, using $\mathbf{H}(t)\leq\mathbf{M}(t)$, one derives
$\operatorname{var}(x_{i,k}(\infty))$ $\leq\frac{1}{\overline{n}}\lim\limits_{t\rightarrow\infty}\sum_{q=0}^{t}\sum_{j=1}^{\overline{n}}\left[\Psi_{t+1,q}^{T}\Psi_{t+1,q}\right]_{jj}\text{var}([\mathbf{v}]_{jk}(q))$
$=\frac{\lambda^{2}}{1-\upsilon^{2}}.$
Then, we establish the accuracy in expectation by [29].
Lemma 4.12.
Under PP-ADRC, if equation (6) holds, and $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, then
$\mathrm{P}\{|x_{i,k}(\infty)-\psi[\overline{\mathbf{x}}_{0}]^{k}|\leq(l_{k}+r_{k})\}$ $\geq 1-\frac{\operatorname{var}(x_{i,k}(\infty))}{(l_{k}+r_{k})^{2}}$ (14)
$\geq 1-\frac{\lambda^{2}/(1-\upsilon^{2})}{(l_{k}+r_{k})^{2}},$
where $l_{k}=\operatorname{min}\{\psi[\overline{\mathbf{x}}_{0}]^{k}-[\overline{\mathbf{x}}_{0}]^{k}_{\operatorname{min}},[\overline{\mathbf{x}}_{0}]^{k}_{\operatorname{max}}-\psi[\overline{\mathbf{x}}_{0}]^{k}\}$, $r_{k}>\max\{0,\lambda\sqrt{\frac{1}{1-\upsilon^{2}}}-l_{k}\}$, and $k=1,2,\ldots,d$.
Proof 4.13.
By Chebyshev’s inequality, we determine the one-dimensional convergence accuracy,
$\mathrm{P}\{|x_{i,k}(\infty)-\psi[\overline{\mathbf{x}}_{0}]^{k}|\leq(l_{k}+r_{k})\}$ $\geq 1-\frac{\operatorname{var}(x_{i,k}(\infty))}{(l_{k}+r_{k})^{2}}$
$\geq 1-\frac{\lambda^{2}/(1-\upsilon^{2})}{(l_{k}+r_{k})^{2}},$
where $r_{k}>\max\{0,\lambda\sqrt{\frac{1}{1-\upsilon^{2}}}-l_{k}\}$ provides a tight lower bound.
Remark 4.14.
Given that $\psi$ is a row vector with elements summing to one, it follows that $\psi[\overline{\mathbf{x}}_{0}]^{k}$ must lie within the range bounded by the maximum and minimum values of $[\overline{\mathbf{x}}_{0}]^{k}$. Consequently, Lemma 4.12 implies that $x_{i,k}(\infty)$ lies within the interval $[\overline{\mathbf{x}}_{0}]^{k}_{\operatorname{max}}+r_{k}$ to $[\overline{\mathbf{x}}_{0}]^{k}_{\operatorname{min}}-r_{k}$ with a probability of at least $1-\frac{\lambda^{2}/(1-\upsilon^{2})}{(l_{k}+r_{k})^{2}}$.
Then, we consider the case that a fraction $(1-\beta)$, with $10<\beta<$ 1, of the state’s dimensions do not need to be protected, i.e., only part of the dimensions are added with noises. The convex hull formed by these dimensions remains unchanged. For the remaining $\beta d$ dimensions, we apply Lemma 4.12 $\beta d$ times. Then, we provide three convex hull definitions.
Definition 4.15.
  1. 1.
    $\mathcal{A}$: Convex hull $\mathcal{A}$ is defined as the convex hull of the initial states of the normal agents.
  2. 2.
    $\mathcal{B}$: Projecting convex hull $\mathcal{A}$ onto $(1-\beta)d$ dimensions without adding noise, we then translate it sequentially along each remaining dimension from its minimum value to its maximum value among all normal agents, forming convex hull $\mathcal{B}$ at the end.
  3. 3.
    $\mathcal{C}$: Convex hull $\mathcal{C}$ is formed by extending the translation range of $\mathcal{B}$, moving from the minimum value minus $r_{k}$ to the maximum value plus $r_{k}$ for the $k$th component of the normal agents’ states.
Then, we can precisely determine the shape of the new convex hull $\mathcal{C}$. We take a $3$-dimensional case as an example, where the coordinates of $6$ points are $(1,0,0)$, $(0,2,0)$, $(-1/4,-1/4,0)$, $(0,0,2)$, $(0,1,0)$, and $(1/4,1/2,0)$. The noise is only added to the z-axis component of the states, ensuring $0\leq\psi[\overline{\mathbf{x}}_{0}]^{3}\leq 2$. The three distinct convex hulls are depicted in Fig. 1. The convex hull $\mathcal{A}$ represents the original convex hull formed by the $6$ points. The convex hulls $\mathcal{B}$ and $\mathcal{C}$ are obtained by translating the part of $\mathcal{A}$ in the $xOy$ plane along the $z$-axis. The convex hull $\mathcal{B}$ translates from 00 to $2$, while $\mathcal{C}$ translates from $-r_{3}$ to $2+r_{3}$. Therefore, utilizing Lemma 4.12, we can calculate that the probability of the final value belonging to the convex hull $\mathcal{C}$ is at least $1-\frac{\lambda^{2}/(1-\upsilon^{2})}{(l_{k}+r_{k})^{2}}$.
(a) $\mathcal{A}$
(b) $\mathcal{A}$ and $\mathcal{B}$
(c) $\mathcal{A}$, $\mathcal{B}$ and $\mathcal{C}$
Figure 1: Comparisons of Three Convex Hulls.
Inspired by [30], we further use Hausdorff distance to measure the distance between the two convex hulls $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, as well as the diameter of a convex hull.
Definition 4.16.
Given two convex hulls $\mathcal{A}_{1}$ and $\mathcal{A}_{2}\subset\mathbb{R}^{d}$, the Hausdorff distance between $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ is defined as
$D_{H}(\mathcal{A}_{1},\mathcal{A}_{2})=\max_{a_{1}\in\partial\mathcal{A}_{1}}\{\min_{a_{2}\in\partial\mathcal{A}_{2}}\{\mathrm{dist}(a_{1},a_{2})\}\},$ (15)
where $\mathrm{dist}(a_{1},a_{2})$ is the Euclidean distance between $a_{1}$ and $a_{2}$, and $\partial\mathcal{A}_{1}$, $\partial\mathcal{A}_{2}$ denote the boundaries of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, respectively.
We denote by $\mu(\mathcal{A}_{1})$ the diameter of $\mathcal{A}_{1}$, i.e.,
$\mu(\mathcal{A}_{1})=\max_{a,b\in\partial\mathcal{A}_{1}}\{\mathrm{dist}(a,b)\}.$ (16)
We utilize both $D_{H}(\mathcal{A}_{1},\mathcal{A}_{2})$ and $\mu(\mathcal{A}_{1})$ to quantify the distance between $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$.
Theorem 4.17.
The convex hull $\mathcal{C}$ satisfies
$D_{H}(\mathcal{A},\mathcal{C})\leq\sqrt{\frac{d}{2}}\;\mu(\mathcal{A})+\sqrt{r_{1}^{2}+r_{2}^{2}+\cdots+r_{\beta d}^{2}},$
where $10<\beta<$ 1, $r_{k}$ is a freely chosen parameter for $k=1,2,\ldots,\beta d$, and $\mathcal{A}$ is the convex hull of the initial states of the normal agents. Under PP-ADRC, if equation (6) holds, $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, and $(1-\beta)$ fraction of the dimensions of the state vector are not added with noises, the probability $\mathrm{P}\{x(\infty)\in\mathcal{C}\}$ that the final value lies within the convex hull $\mathcal{C}$ satisfies
$\mathrm{P}\{x(\infty)\in\mathcal{C}\}\geq\prod\limits_{k=1}^{\beta d}\left[1-\frac{\lambda^{2}/(1-\upsilon^{2})}{(l_{k}+r_{k})^{2}}\right].$ (17)
Proof 4.18.
We characterize the upper bound of $D_{H}(\mathcal{A},\mathcal{C})$ by splitting it into $D_{H}(\mathcal{A},\mathcal{B})$ and $D_{H}(\mathcal{B},\mathcal{C})$. Thus, the proof is divided into two parts.
$D_{H}(\mathcal{A},\mathcal{B})$: Let us discuss a special case as presented in [30]. We define a new convex hull $\mathcal{D}=\mathcal{A}_{1}\times\mathcal{A}_{2}\times\ldots\times\mathcal{A}_{d}$, where $\mathcal{A}_{k}=\operatorname{conv}(x_{1,k}(0),x_{2,k}(0),\ldots,x_{\overline{n},k}(0))$, $k=1,2,\ldots,d$. The example in $\mathbb{R}^{3}$ is shown in Fig. 2. Apparently, $D_{H}(\mathcal{A},\mathcal{D})$ is bound to greater than or equal to $D_{H}(\mathcal{A},\mathcal{B})$. Moreover, it is also obtained that $D_{H}(\mathcal{A},\mathcal{D})\leq\sqrt{\frac{d}{2}}\;\mu(\mathcal{A})$. Therefore, we obtain $D_{H}(\mathcal{A},\mathcal{B})\leq D_{H}(\mathcal{A},\mathcal{D})\leq\sqrt{\frac{d}{2}}\;\mu(\mathcal{A})$.
(a) $\mathcal{A}$ and $\mathcal{D}$
(b) $\mathcal{A}$, $\mathcal{B}$, and $\mathcal{D}$
Figure 2: Hyperrectangle in $\mathbb{R}^{3}$.
$D_{H}(\mathcal{B},\mathcal{C}):$ we consider a surface of $\mathcal{B}$ and denote its vertices by $\mathcal{B}_{1},\ldots,\mathcal{B}_{\kappa}$. We then characterize the nearest surface of $\mathcal{C}$ parallel to the surface we just mentioned, and the corresponding vertices are $\mathcal{C}_{1},\ldots,\mathcal{C}_{\kappa}$. by recalling the definition of the Hausdorff distance, it must correspond to the distance between points on two surfaces. We denote by $b$ and $c$ any two points on the surface of $\mathcal{B}$ and $\mathcal{C}$, respectively. We can obtain $b=b_{1}\mathcal{B}_{1}+\ldots+b_{\kappa}\mathcal{B}_{\kappa}$ and $c=c_{1}\mathcal{C}_{1}+\ldots+c_{\kappa}\mathcal{C}_{\kappa}$, where $b_{1}+\ldots+b_{\kappa}=c_{1}+\ldots+c_{\kappa}=1$, and $b_{i}\geq 0,c_{i}\geq 0,\ i=1,2,\ldots,\kappa$. Moreover, from Definition 4.15, we can obtain $\mathcal{B}_{i}-\mathcal{C}_{i}=[\pm r_{1},\pm r_{2},\ldots,\pm r_{\beta d}]^{\top},i=1,2,\ldots,\kappa$, where the sign depends on the position of the surface. Hence, the distance $\mathrm{dist}(b,c)$ between $b$ and $c$ satisfies
$\mathrm{dist}(b,c)$ $=\sqrt{\sum\limits_{i=1}^{\kappa}||b_{i}\mathcal{B}_{i}-c_{i}\mathcal{C}_{i}||_{2}^{2}}$
$=\sqrt{\sum\limits_{i=1}^{\kappa}\left[(b_{i}-c_{i})^{2}||\mathcal{C}_{i}||_{2}^{2}+b_{i}^{2}(r_{1}^{2}+\ldots+r_{\beta d}^{2})\right]}.$
First, we assume $b_{i}$ is a constant and then find the lower bound of $\mathrm{dist}(b,c)$. Apparently when $c_{i}=b_{i}$, $\mathrm{dist}(b,c)$ is lower bounded, which is $\min\limits_{c_{i}}(\mathrm{dist}(b,c))=\sqrt{\sum\limits_{i=1}^{\kappa}\left[b_{i}^{2}(r_{1}^{2}+\ldots+r_{\beta d}^{2})\right]}$. Then, the problem becomes finding the upper bound of $\min\limits_{c_{i}}(\mathrm{dist}(b,c))$, which is denoted by $\max\limits_{b_{i}}\min\limits_{c_{i}}(\mathrm{dist}(b,c))$. Given that $b_{1}+\ldots+b_{\kappa}=1$ and $b_{1}^{2}+\ldots+b_{\kappa}^{2}\leq(b_{1}+\ldots+b_{\kappa})^{2}$, we can know that the upper bound of $b_{1}^{2}+\ldots+b_{\kappa}^{2}$ is $1$ when $b_{i_{0}}=1,\ i_{0}\in{1,2,\ldots,\kappa}\ \text{ and }\ b_{i}=0,\ i\neq i_{0}$. Therefore, we can obtain $\max\limits_{b_{i}}\min\limits_{c_{i}}(\mathrm{dist}(b,c))=\sqrt{r_{1}^{2}+\ldots+r_{\beta d}^{2}}$. According to the definition, $\max\limits_{b_{i}}\min\limits_{c_{i}}(\mathrm{dist}(b,c))$ is exactly the Hausdorff distance $D_{H}(\mathcal{B},\mathcal{C})$ between $\mathcal{B}$ and $\mathcal{C}$. So we can obtain
$D_{H}(\mathcal{A},\mathcal{C})$ $\leq D_{H}(\mathcal{A},\mathcal{B})+D_{H}(\mathcal{B},\mathcal{C})$
$\leq\sqrt{\frac{d}{2}}\;\mu(\mathcal{A})+\sqrt{r_{1}^{2}+r_{2}^{2}+\cdots+r_{\beta d}^{2}}.$
The corresponding probability satisfies
$\mathrm{P}\{x(\infty)\in\mathcal{C}\}$ $\geq\prod\limits_{k=1}^{\beta d}\left[1-\frac{\operatorname{var}(x_{i,k}(\infty)}{(l_{k}+r_{k})^{2}}\right]$
$\geq\prod\limits_{k=1}^{\beta d}\left[1-\frac{\lambda^{2}/(1-\upsilon^{2})}{(l_{k}+r_{k})^{2}}\right].$
Remark 4.19.
The one-dimensional results can yield an analytical solution because there is no need to consider the correlations between dimensions, making the analysis easier. The probability of convergence accuracy is influenced by several parameters. Specifically, if the Hausdorff distance between the convex hulls increases, the probability decreases. Similarly, an increase in the noise parameters $\lambda$ and $\upsilon$, which corresponds to an increase in variance, also leads to a decrease in probability.
5 Privacy Analysis
In this section, we analyze the $\rho-$concentrated geo-privacy of PP-ADRC and compare it with $(\varepsilon,\delta)-$differential privacy. At last, we make a detailed discussion on the relationship between convergence accuracy and privacy.
5.1 $\rho-$Concentrated Geo-Privacy
We first investigate the $\rho-$CGP properties of PP-ADRC. As defined in Definition 8, in the resilient vector consensus, the input is the initial states of normal agents $\overline{\mathbf{x}}_{0}$ and the output is the infinite sequence of noised states $\mathbf{Y}=\{\overline{\mathbf{y}}(t)\}_{t=0}^{\infty}$.
Theorem 5.20.
Under PP-ADRC, if equation (6) holds, and $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, then we achieve $\rho-$CGP, where $\rho=\bigg{[}\frac{n\upsilon^{2}}{2\lambda^{2}[\upsilon^{2}-(1-\gamma_{l})^{2}]}\bigg{]}$.
Proof 5.21.
Here, we present a general proof overview: Initially, we identify a variable that equalizes the outputs under two initial conditions following those in [29, 20]. Subsequently, we apply a variable transformation to standardize the integral domain and compute the Rényi divergence.
For any pair of initial conditions $\overline{\mathbf{x}}_{0}$ and $\overline{\mathbf{x}}_{0}^{\prime}$ and an arbitrary set $\mathcal{O}\in\mathcal{B}((\mathbb{R}^{\overline{n}\times d})^{\mathbb{N}})$ of output, we consider the corresponding input domain. For any $k\geq 0$, let the domain $R_{k}=\{\mathbf{N}_{k}\in\Omega_{k}|\mathbf{Y}_{k,\overline{\mathbf{x}}_{0}}(\mathbf{N}_{k})\in\mathcal{O}_{k}\},$ where $\Omega_{k}=(\mathbb{R}^{\overline{n}\times d})^{k+1}$ is the sample space until time $k$ and $\mathcal{O}_{k}$ is the output set obtained by truncating the elements of $\mathcal{O}$ to finite subsequences of length $k+1$. We have the same definition for $\overline{\mathbf{x}}_{0}^{\prime}$, which is $R_{k}^{\prime}=\{\mathbf{N}_{k}\in\Omega_{k}|\mathbf{Y}_{k,\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N}_{k})\in\mathcal{O}_{k}\}$. Based on the continuity of probability [31], the probability can be written as
$\mathrm{P}\{\mathbf{N}\in\Omega|\mathbf{Y}_{\overline{\mathbf{x}}_{0}}(\mathbf{N})\in\mathcal{O}\}=\lim_{k\to\infty}\int_{R_{k}}\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k})\mathrm{d}\mathbf{N}_{k}$ (18)
and
$\mathrm{P}\{\mathbf{N}\in\Omega|\mathbf{Y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N})\in\mathcal{O}\}=\lim_{k\to\infty}\int_{R_{k}^{\prime}}\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k}^{\prime})\mathrm{d}\mathbf{N}_{k}^{\prime},$ (19)
where $\mathsf{f}_{\overline{n}(k+1)}$ represents the pdf of a joint multivariate Gaussian distribution in $\overline{n}(k+1)$ dimensions, defined as
$\begin{split}\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k})&=\prod\limits_{h=0}^{k}\prod\limits_{i=1}^{\overline{n}}G(\eta_{i}(h);\Sigma(h)).\end{split}$ (20)
Therefore, we can determine the distribution of the output of the first execution, $\mathbf{y}_{\overline{\mathbf{x}}_{0}}(\mathbf{N})=\lim\limits_{k\to\infty}\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k})$, and the output of the second execution, $\mathbf{y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N})=\lim\limits_{k\to\infty}\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k}^{\prime})$. The next step is to calculate the Rényi divergence of these two distributions, but the integral variables are different, making it unrealizable for us to make a direct computation. Therefore, we need to find the relationship between $\mathbf{N}$ and $\mathbf{N}^{\prime}$ so that the integral variables and domains of $\mathbf{y}_{\overline{\mathbf{x}}_{0}}(\mathbf{N})$ and $\mathbf{y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N}^{\prime})$ are the same, differing only in the integral expression. Given that $\overline{\mathbf{x}}_{0}$ and $\overline{\mathbf{x}}_{0}^{\prime}$ are two arbitrary initial conditions, we assume that $x_{i}^{\prime}(0)=x_{i}(0)+\delta_{i}$, $i=1,2,\ldots,\overline{n}$, where $\delta_{i}\in\mathbb{R}^{d}$ and $\|\delta_{i}\|_{2}\leq\mathrm{dist}(\overline{\mathbf{x}}_{0},\overline{\mathbf{x}}_{0}^{\prime})$. Then, for any $\mathbf{N}_{k}\in R_{k}$, we define $\mathbf{N}_{k}^{\prime}$ by
$\eta_{i}^{\prime}(h)=\begin{cases}\eta_{i}(h)-\prod_{t=0}^{h-1}\left[1-\gamma_{i}(t)\right]\delta_{i}&\text{if }h>0.\\ \eta_{i}(h)-\delta_{i}&\text{if }h=0,\end{cases}$ (21)
for $i=1,2,\cdots,\overline{n}$.
Then, we need to show that based on the above definitions, the outputs $\mathbf{Y}_{k,\overline{\mathbf{x}}_{0}}(\mathbf{N}_{k})$ and $\mathbf{Y}_{k,\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N}_{k}^{\prime})$ have the same distribution. For the normal agents, we use the method of mathematical induction. From equation (3), we can derive $y_{i}^{\prime}(0)=x_{i}^{\prime}(0)+\eta_{i}^{\prime}(0)=x_{i}(0)+\delta_{i}+\eta_{i}(0)-\delta_{i}=y_{i}(0),i=1,2,\cdots,\overline{n}$. We can conclude that for every normal agent $i$, the calculated centerpoints $s_{i}(0)$ and $s_{i}^{\prime}(0)$ are the same. According to equation (4), we can derive $x_{i}^{\prime}(1)=x_{i}(1)+(1-\gamma_{i}(0))\delta_{i}$. By induction, we can easily obtain that $x_{i}^{\prime}(h)=x_{i}(h)+\prod_{t=0}^{h-1}\left[1-\gamma_{i}(t)\right]\delta_{i},h=1,2,\cdots,k$, followed by $y_{i}(h)=y_{i}^{\prime}(h)$. It means that $\mathbf{Y}_{k,\overline{\mathbf{x}}_{0}}(\mathbf{N}_{k})=\mathbf{Y}_{k,\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N}_{k}^{\prime})$, indicating that $\mathbf{N}_{k}^{\prime}\in R_{k}^{\prime}$ and we successfully build a bijective correspondence. For any $\mathbf{N}_{k}^{\prime}\in R_{k}^{\prime}$, there exists $(\mathbf{N}_{k},\Delta\mathbf{N}_{k})\in R_{k}\times(\mathbb{R}^{\overline{n}\times d})^{k+1}$ such that $\mathbf{N}_{k}^{\prime}=\mathbf{N}_{k}+\Delta\mathbf{N}_{k}$. Therefore, we derive the integral by changing variables $\mathrm{P}\{\mathbf{N}\in\Omega|\mathbf{Y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N})\in\mathcal{O}\}=\lim\limits_{k\to\infty}\int_{R_{k}}\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k}+\Delta\mathbf{N}_{k})\mathrm{d}\mathbf{N}_{k},$ indicating that the distribution $\mathbf{y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N})$ can be expressed as $\lim\limits_{k\rightarrow\infty}\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k}+\Delta\mathbf{N}_{k})$.
With all the preparations above, we can further analyze the Rényi divergence of $\mathbf{y}_{\overline{\mathbf{x}}_{0}}(\mathbf{N})$ and $\mathbf{y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N})$ as equation (22).
$D_{\alpha}(\mathbf{y}_{\overline{\mathbf{x}}_{0}}(\mathbf{N})\|\mathbf{y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N}))$ (22)
$=\lim_{k\to\infty}\frac{1}{\alpha-1}\mathrm{log}\bigg{[}\int[\mathbf{y}_{\overline{\mathbf{x}}_{0}}(\mathbf{N}_{k})]^{\alpha}[\mathbf{y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N}_{k})]^{1-\alpha}\text{d}\mathbf{N}_{k}\bigg{]}$
$=\lim_{k\to\infty}\frac{1}{\alpha-1}\mathrm{log}\bigg{[}\int_{R_{k}}[\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k})]^{\alpha}[\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k}+\Delta\mathbf{N}_{k})]^{1-\alpha}\mathrm{d}\mathbf{N}_{k}\bigg{]}$
Then, we need to calculate detailed expressions of $[\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k})]^{\alpha}[\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k}+\Delta\mathbf{N}_{k})]^{1-\alpha}$.
$[\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k})]^{\alpha}[\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k}+\Delta\mathbf{N}_{k})]^{1-\alpha}$ (23)
$=\bigg{[}\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k})\bigg{]}^{\alpha}\times\bigg{[}\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k}+\Delta\mathbf{N}_{k})\bigg{]}^{1-\alpha}$
$=\prod\limits_{h=0}^{k}\bigg{[}\prod\limits_{i=1}^{\overline{n}}G(\eta_{i}(h);\Sigma(h))\bigg{]}^{\alpha}\times\bigg{[}\prod\limits_{i=1}^{\overline{n}}G(\eta_{i}(h)+\Delta\eta_{i}(h);\Sigma(h))\bigg{]}^{1-\alpha}$
$=\prod\limits_{h=0}^{k}\prod\limits_{i=1}^{\overline{n}}\bigg{[}\frac{1}{\sqrt{(2\pi\sigma^{2}(h))^{d}}}\mathrm{exp}\left(-\frac{\alpha||\eta_{i}(h)||^{2}_{2}}{2\sigma^{2}(h)}\right)\cdot\mathrm{exp}\left(-\frac{(1-\alpha)||\eta_{i}(h)+\Delta\eta_{i}(h)||^{2}_{2}}{2\sigma^{2}(h)}\right)\bigg{]}$
$=\prod\limits_{h=0}^{k}\prod\limits_{i=1}^{\overline{n}}\bigg{[}\frac{1}{\sqrt{(2\pi\sigma^{2}(h))^{d}}}\mathrm{exp}\left(-\frac{||\eta_{i}(h)-(1-\alpha)\Delta\eta_{i}(h)||^{2}_{2}-||(1-\alpha)\Delta\eta_{i}(h)||^{2}_{2}+(1-\alpha)||\Delta\eta_{i}(h)||^{2}_{2}}{2\sigma^{2}(h)}\right)\bigg{]}$
Substituting equation (23) into (22) , we can obtain the Rényi divergence between the two distributions.
$D_{\alpha}(\mathbf{y}_{\overline{\mathbf{x}}_{0}}(\mathbf{N})\|\mathbf{y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N}))$ (24)
$=\lim_{k\to\infty}\frac{1}{\alpha-1}\mathrm{log}\bigg{[}\int_{R_{k}}\prod\limits_{h=0}^{k}\prod\limits_{i=1}^{\overline{n}}\bigg{[}\frac{1}{\sqrt{(2\pi\sigma^{2}(h))^{d}}}\mathrm{exp}\left(-\frac{||\eta_{i}(h)-(1-\alpha)\Delta\eta_{i}(h)||^{2}_{2}-||(1-\alpha)\Delta\eta_{i}(h)||^{2}_{2}+(1-\alpha)||\Delta\eta_{i}(h)||^{2}_{2}}{2\sigma^{2}(h)}\right)\bigg{]}$
$\mathrm{d}\eta_{1}(h)\cdots\eta_{\overline{n}}(h)$
$=\lim_{k\to\infty}\frac{1}{\alpha-1}\mathrm{log}\prod\limits_{h=0}^{k}\{\prod\limits_{i=1}^{\overline{n}}\underset{\eta_{i}(h){\sim}\mathcal{MVN}((1-\alpha)\Delta\eta_{i}(h),\sigma^{2}(h)I_{d})}{\mathbb{E}}\bigg{[}\mathrm{exp}\left(-\frac{-||(1-\alpha)\Delta\eta_{i}(h)||^{2}_{2}+(1-\alpha)||\Delta\eta_{i}(h)||^{2}_{2}}{2\sigma^{2}(h)}\right)\bigg{]}\}$
$=\lim_{k\to\infty}\sum\limits_{h=0}^{k}\bigg{[}\sum\limits_{i=1}^{\overline{n}}\frac{\alpha||\Delta\eta_{i}(h)||^{2}_{2}}{2\lambda^{2}\upsilon^{2h}}\bigg{]}\leq\bigg{[}\frac{\alpha n}{2\lambda^{2}}+\lim_{k\to\infty}\sum\limits_{h=1}^{k}\frac{\alpha n(1-\gamma_{l})^{2h}}{2\lambda^{2}\upsilon^{2h}}\bigg{]}\mathrm{dist}(\overline{\mathbf{x}}_{0},\overline{\mathbf{x}}_{0}^{\prime})^{2}$
$\leq\alpha\bigg{[}\frac{n\upsilon^{2}}{2\lambda^{2}[\upsilon^{2}-(1-\gamma_{l})^{2}]}\bigg{]}\mathrm{dist}(\overline{\mathbf{x}}_{0},\overline{\mathbf{x}}_{0}^{\prime})^{2},$
where $\mathcal{MVN}((1-\alpha)\Delta\eta_{i}(h),\sigma^{2}(h)I_{d})$ represents a multivariate Gaussian distribution with mean $(1-\alpha)\Delta\eta_{i}(h)$ and covariance matrix $\sigma^{2}(h)I_{d}$. Consequently, we have found the upper bound of the Rényi divergence, thereby completing the proof.
Remark 5.22.
We analyze the privacy of infinite sequences as outputs for two reasons. First, due to the noise adding, convergence is non-deterministic and can theoretically only be achieved as the iteration approaches infinity, when the noise diminishes to a negligible level. Secondly, since the output at each iteration can potentially leak the initial states to varying degrees, we apply a composition conclusion to analyze the upper bound of privacy leakage. It is important to note that we only consider the states of normal agents and neglect the states of the faulty agents, as we are concerned with the state evolution of normal agents under PP-ADRC. Normal agents may leak privacy when sending state messages; thus, this issue is independent of the status and transmitted information of the faulty agents. However, the states of normal agents can be influenced by the states of faulty agents. Moreover, given variables $\lambda$ and $\upsilon$, we plot the surface of $\rho$ concerning them, as shown in Figure 3. It can be observed that $\rho$ monotonically increases with both $\lambda$ and $\upsilon$.
Figure 3: Change trend of $\rho$
5.2 Comparison with $(\varepsilon,\delta)-$Differential Privacy
We first give the definition of $(\varepsilon,\delta)-$DP in the resilient vector consensus.
Definition 5.23.
Given $\ell\in\mathbb{R}\geq 0$, the initial states of normal agents $\overline{\mathbf{x}}_{0}$ and $\overline{\mathbf{x}}_{0}^{\prime}$ are $\ell$-neighboring if, for some $i_{0}\in\overline{\mathcal{V}},||[\overline{\mathbf{x}}_{0}]_{i_{0}}-[\overline{\mathbf{x}}_{0}^{\prime}]_{i_{0}}||_{2}\leq\ell$ and $[\overline{\mathbf{x}}_{0}]_{i}=[\overline{\mathbf{x}}_{0}^{\prime}]_{i}$ if $i\neq i_{0}$ and $i\in\overline{\mathcal{V}}$. Given $\ell$, $\varepsilon$ and $\delta\in\mathbb{R}\geq 0$, the system achieves $(\varepsilon,\delta)$-differentially private if, for any pair of $\overline{\mathbf{x}}_{0}$ and $\overline{\mathbf{x}}_{0}^{\prime}$ of $\ell-$neighboring initial states and any set $\mathcal{O}\in\mathcal{B}((\mathbb{R}^{\overline{n}\times d})^{\mathbb{N}})$,
$\mathrm{P}\{\mathbf{N}\in\Omega|\mathbf{Y}_{\overline{\mathbf{x}}_{0}}(\mathbf{N})\in\mathcal{O}\}\leq e^{\varepsilon}\mathrm{P}\{\mathbf{N}\in\Omega|\mathbf{Y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N})\in\mathcal{O}\}+\delta.$ (25)
Remark 5.24.
The definition of neighboring initial states here differs from those in existing papers [21, 20, 29], where the $1$-norm is used. In those papers, differential privacy is $\epsilon$-DP with Laplace noise, whereas we use $(\epsilon,\delta)$-DP with Gaussian noise. Therefore, the neighboring initial states are defined using the $2$-norm.
As we have talked before, the output of the system is an infinite sequence. The output at each time can be considered as a query. Therefore, the final result is the infinite composition of each single query at time $h$ satisfying $(\varepsilon(h),\delta(h))-$DP. We now give the detail of $\varepsilon(h)$ and $\delta(h)$ according to Theorem A.1. in [32].
Lemma 5.25.
For $h=0,1,2,\ldots$, the output $\overline{\mathbf{y}}(h)$ at each time $h$ can be considered as a query, satisfying $(\varepsilon(h),\delta(h))-$DP. Given $\delta(h)\in\mathbb{R}>0$, $\varepsilon(h)$ holds that $\varepsilon(h)>\frac{\sqrt{2\mathrm{log}\frac{1.25}{\delta(h)}}\ell(1-\gamma_{l})^{h}}{\lambda\upsilon^{h}}$.
Regarding the composition, we can simply sum $\varepsilon(h)$ and $\delta(h)$ from $h=0$ to $\infty$. This results in rapid growth of privacy loss and $\delta$, indicating a poor outcome. However, computing the tightest possible privacy guarantee for such a composition is #P-hard [33]. Then, we can make a detailed comparison between these two definitions below:
  • CGP better meets the privacy preservation needs. In resilient vector consensus, our privacy preservation target is the initial states of all normal agents, which may be leaked during communication, rather than privacy leakage caused by the alteration or addition/deletion of an initial state in a data center. Therefore, the definition of $\ell$-neighboring states in DP does not fully meet our needs, as it should consider any pair of initial states. Consequently, CGP describes the degree of output differences for any pair of inputs, making it more suitable for protecting the initial state of each normal agent under resilient vector consensus. It also uses the distance between initial states as a privacy protection parameter, thereby greatly improving the system’s utility.
  • CGP admits better advanced composition. In advanced composition, privacy loss is linear with the square root of the number of queries. Although $(\varepsilon,\delta)-$DP admits advanced composition, the computation process is quite complicated and would make $\delta$ grow linearly, increasing the risk of privacy leakage. However, the advanced composition of CGP is quite straightforward: it involves simply summing up the $\rho$ values without considering any additional parameters. In resilient vector consensus, we consider the noisy states sent by the agents from time zero to infinity as outputs. Each output at every iteration is considered a potential privacy leakage query. The final result we obtain is essentially the outcome of an infinite number of queries. Using CGP results in less privacy loss and easier computation compared to $(\varepsilon,\delta)-$DP, particularly in scenarios with a large number of queries, as it provides a stricter and more accurate upper bound on privacy preservation.
5.3 Trade-off between Privacy and Accuracy
In this part, we discuss the tradeoff between privacy and accuracy. As we analyze the convergence accuracy from two perspectives, i.e., distribution of the final value and convex hull change, we also detail the tradeoff from these two perspectives. From Theorem 5.20, the parameters that determine $\rho$ are provided. The parameters that we can freely choose are the noise parameters $\lambda$ and $\upsilon$. The larger $\lambda$ and $\upsilon$ are, the stronger the privacy preservation is. Therefore, we analyze the relationship between accuracy and privacy by changing $\lambda$ and $\upsilon$.
  • Distribution of the final value: The Mahalanobis distance is a normalized coefficient, meaning its value is fixed artificially and remains uninfluenced by variance or covariance. However, the volume of the hyperspheroid formed at an equal Mahalanobis distance can be influenced by various parameters. Therefore, the metric that accurately represents precision should be the volume of the hyperspheroid: the larger the volume, the lower the precision. As previously analyzed, increasing the amplitude parameters $\lambda$ and $\upsilon$ of the noise increases the volume of the hyperspheroid, thereby reducing convergence accuracy and decreasing $\rho$, which corresponds to enhanced privacy protection.
  • Convex hull change: We use the Hausdorff distance to characterize the distance between convex hulls before and after noise adding. Similar to the Mahalanobis distance, the Hausdorff distance is also specified artificially. However, the probability of convergence within the convex hull is influenced by the noise parameters. The larger the noise amplitude parameters $\lambda$ and $\upsilon$, the lower the probability that the new convex hull will converge within a given Hausdorff distance. This results in a decrease in convergence accuracy but simultaneously leads to a reduction in $\rho$, thereby enhancing the level of privacy protection.
6 Simulations
In this section, we conduct extensive simulations of a multi-robot system performing PP-ADRC both in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$ to illustrate our theoretical results.
6.1 $2-$dimensional case
In this scenario, we deploy $10$ robots with a time-varying directed network, consisting of $2$ faulty robots and $8$ normal robots. The faulty robot sends its position to every normal robot. Each normal robot’s in-neighborhood is selected randomly at each iteration, while ensuring the conditions of resilience. The randomly generated initial states are depicted in Fig. 4(a), where blue points represent normal robots’ states and the red ones are the states of faulty robots. The green polygon represents the convex hull of normal robots. The faulty robots oscillate following the rule in [17]. For parameters, we set $\alpha_{i}(t)=0.8,\ \lambda=2.0$, and $\upsilon=0.75$.
(a) Initial states
(b) Final value
Figure 4: $2$-dimensional case
We execute the algorithm $1000$ times, each with $1000$ iterations, and obtain the result in Fig. 4(b), where blue points represent the final values each time, and the green point $(-0.011,0.210)$ is the sample expectation of blue points. We observe that the expectation falls in the convex hull of the initial states of normal robots.
Then, we only add noise to the second dimension to illustrate Theorem 4.17. We repeated the process $1000$ times, with $1000$ iterations for three different cases of noise parameters (case 1): $\lambda=2.0,\upsilon=0.75$, case 2): $\lambda=2.5,\upsilon=0.75$, and case 3): $\lambda=2.5,\upsilon=0.85$) and obtain the convergence results shown in Fig. 5. The yellow convex hull represents $\mathcal{C}$, with $r_{2}=0.3$. The variances of the second dimension in order are $0.138,0.237,$ and $0.336$, respectively. In Fig. 5(d), the vertical axis represents the probability of the final value falling within the convex hull and the horizontal axis represents $r_{2}$. There are two sets of lines, i.e., one obtained through simulation probabilities $\mathrm{P}^{\rm Si}$ and the other $\mathrm{P}^{\rm Th}$ derived theoretically from Theorem 4.17. It is observed that in all three cases, the simulation results $\mathrm{P}^{\rm Si}$ are greater than the theoretical ones $\mathrm{P}^{\rm Th}$. When the distance satisifies $r_{2}=0.3$, we have $\delta(\mathcal{A},\mathcal{C})=0.919$ and $\mu(\mathcal{A})=1.77$, which illustrates our theoretical result $\delta(\mathcal{A},\mathcal{C})\leq\sqrt{\frac{d}{2}}\mu(\mathcal{A})+r_{2}$.
(a) $\lambda=2.0,\upsilon=0.75$
(b) $\lambda=2.5,\upsilon=0.75$
(c) $\lambda=2.5,\upsilon=0.85$
(d) Probability
Figure 5: $2-$dimensional Hausdorff Distance Results
At last, we illustrate the final value distribution using Mahalanobis distance. For the $2-$dimensional case, We also repeated the process $1000$ times, with $1000$ iterations for three different cases of noise parameters as below (case 1): $\lambda=2.0,\upsilon=0.75$, case 2): $\lambda=2.5,\upsilon=0.75$, and case 3): $\lambda=2.0,\upsilon=0.65$), and obtain the convergence result shown in Fig. 6. Considering that the set of points equidistant by Mahalanobis distance forms an ellipse in two dimensions, we have drawn $4$ red ellipses under different parameters of $\chi$ for each figure, which are $2,3,4,5$ in order. It can be observed that $\mathrm{P}^{\mathrm{Si}}$ are greater than the theoretical one $\mathrm{P}^{\mathrm{Th}}$, illustrating our results in Theorem 4.6.
(a) $\lambda=2.0,\upsilon=0.75$
(b) $\lambda=2.5,\upsilon=0.75$
(c) $\lambda=2.0,\upsilon=0.65$
(d) Probability
Figure 6: $2-$dimensional Mahalabinos Distance Results
6.2 $3-$dimensional case
For the $3$-dimensional case, according to [34], although we can find a centerpoint with depth of $\frac{n}{d+1}=\frac{n}{4}$, but the running time is $O(n^{2})$, make the algorithm inefficient. Therefore, Har-Peled et al. proposed a compromised algorithm that achieves a depth of $\frac{n}{6}$ for the centerpoint with a runtime of $O(n\operatorname{log}n)$, which is used in our simulations here. We consider multi-robots rendezvous, where there are $12$ robots with a time-varying directed network, consisting of $2$ faulty robots and $10$ normal robots. Each normal robot interacts with one randomly chosen faulty robot and other normal robots, where the sufficient conditions for resilience are always guaranteed. The initial states are shown in Fig. 7(a), where the gray polyhedron represents the convex hull formed by normal robots. The faulty robots oscillate, with $\alpha_{i}(t)=0.8$, $\lambda=2.0$, and $\upsilon=0.75$, the same as the 2-dimensional case. We run the algorithm $1000$ times, each with $1000$ iterations, and obtain the result in Fig. 7(b). The blue points still represent final values each time, and the green point $(0.051,0.250,-0.312)$ is the sample expectation of final values. We can also observe that the expectation falls in the convex hull of the initial states of normal robots.
(a) Initial states
(b) Final value
Figure 7: $3-$dimensional case
Then, similar to the $2-$dimensional case, we only add noise to the first dimension to illustrate the convex hull change using Hausdorff distance. There are three different cases of noise parameters (case 1): $\lambda=3.0,\upsilon=0.80$, case 2): $\lambda=3.0,\upsilon=0.85$, and case 3): $\lambda=3.5,\upsilon=0.85$). The results shown in Fig. 8 illustrate our results in Theorem 4.17. The yellow convex hull represents $\mathcal{C}$, with $r_{1}=0.3$. The variances of the first dimension in order are $0.307$, $0.520$, and $0.737$, respectively. When the distance satisifies $r_{1}=0.3$, we have $\delta(\mathcal{A},\mathcal{C})=1.255$ and $\mu(\mathcal{A})=2.75$, which illustrates our theoretical result $\delta(\mathcal{A},\mathcal{C})\leq\sqrt{\frac{d}{2}}\mu(\mathcal{A})+r_{1}$.
(a) $\lambda=3.0,\upsilon=0.80$
(b) $\lambda=3.0,\upsilon=0.85$
(c) $\lambda=3.5,\upsilon=0.85$
(d) Probability
Figure 8: $3-$dimensional Hausdorff Distance Results
Finally, we conduct the simulation to illustrate the final value distribution using Mahalanobis distance in the $3-$dimensional case shown in Fig. 9. The noise parameters are the same as the $2-$dimensional case, while for the gray hyperspheroid shown in the figure, we choose $\chi=3$. From Fig. 9(d), we observe that $\mathrm{P}^{\mathrm{Si}}$ are greater than the theoretical one $\mathrm{P}^{\mathrm{Th}}$, demonstrating out theoretical results in Theorem 4.6.
(a) $\lambda=2.0,\upsilon=0.75$
(b) $\lambda=2.5,\upsilon=0.75$
(c) $\lambda=2.0,\upsilon=0.65$
(d) Probability
Figure 9: $3-$dimensional Mahalabinos Distance Results
7 Conclusion
In this paper, we investigated private resilient vector consensus by adding Gaussian noise for a multi-agent system where there might be faulty agents. We first showed the resilient vector consensus for the normal agents in expectation. For the convergence accuracy, We proposed to analyze from the perspective of Mahalanobis distance between the final value and its expectation and Hausdorff distance between the two convex hulls with and without noise, respectively. Then we proved the $\rho-$concentrated geo-privacy for their initial states, and compare it with $(\varepsilon,\delta)-$differential privacy. At last, we demonstrated our result through numerical simulations. In the future, we will figure out the detailed distribution of the final value instead of using Chebyshev’s inequality. Moreover, it is desirable to compare the performance of algorithms with Laplace noise and Gaussian noise thoroughly.